How many moles of watr

gas?

are needed to produce 5.00 x1022 molecules of hydrogen


1
Expert's answer
2022-04-06T05:01:04-0400

Solution:

hydrogen gas = H2

water = H2O


One mole of any substance contains 6.02×1023 atoms/molecules.

Therefore,

Moles of H2 = (5.00×1022 molecules) × (1 mol / 6.02×1023 molecules) = 0.083 mol


Balanced chemical equation:

2H2O(g) → 2H2(g) + O2(g)

According to stoichiometry:

2 mol of H2O produces 2 mol of H2

X mol of H2O produces 0.083 mol of H2

Thus:

Moles of H2O = X = (0.083 mol H2) × (2 mol H2O / 2 mol H2) = 0.083 mol H2O

Moles of H2O = 0.083 mol


Answer: 0.083 moles of water (H2O) are needed